if all the conditions were met got hardy-weinberg equilibruim in a population of which the dominant alllr represented .64 (64%) of the population. what would be yhe expected outcome of the F1 generation? (p + q) ^2
Answer #1

According to hardy-weinberg concept,

( p+q )2 = p2 + q2 + 2pq

where, p = frequency of dominant allele

q = frequency of recessive allele

Now, given is p2 = 0.64

therefore, p = 0.8

and, q = 0.2

Putting these values in the equation,

we get,

( p + q )2 = ( 0.8)2 + (0.2)2 + 2(0.8)(0.2)

= 0.64 + 0.04 + 0.32

Therefore, In F1 generation,

Homozygous Dominant = 0.64 or 64%

Homozygous Recessive = 0.04 or 4%

Heterozygous = 0.36 or 36%